• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 23:21:45 on May 2,2012

    Entry number 381562

    keyword=Septum Magnets for Tuning

    We have set the septum magnets as Jixie suggested for preparation for
    beam tuning. The regulation has been turned off for both magnets for
    now.

    LHRS: septum current = 362 A
    RHRS: septum current = 475 A

    Once beam is tuned and we are ready for optics, then the regulation
    should be turned back on. The expected currents are

    LHRS P0 = 2.251, left septum current = 725.4 A
    RHRS P0 = 2.161, right septum current = 920 A

    Please note Jack's warning about running at currents at 925 Amps or
    greater for the right septum magnet.


    FIGURE 1